# Axial Seamount Eruption Forecasting Experiment

### v1

@@ Abstract
* Correspondence: qinghua.lei@geo.uu.seAuthor contributions. Conceptualisation: Q.L. and D.S.; Experiment design: Q.L., D.S., and W.W.C.; Code development: Q.L., D.S., and P.S.; Data acquisition: W.W.C., S.N., W.S.D.W., and M.Z.; Forecast analysis: Q.L. and D.S.; Document writing: Q.L., D.S., W.W.C., and M.Z.; Scientific discussion: all authors.
@@ 2 Experimental Design
  • Forecast release schedule. We will issue forecasts on a regular schedule announced in advance on the website of the Geohazards Crisis Observatory (www.geohazards-observatory.com). The first release is planned for 8 November 2025. Thereafter, we aim to provide forecast updates at least once a month, or more frequently as conditions warrant.
@@
  • Digital fingerprinting. The forecast document will not be publicly released immediately. Instead, we will compute its digital fingerprint using a modern cryptographic hash function (SHA-256) under a GNU/Linux environment, in compliance with current recommendations by the National Institute of Standards and Technology. This produces a unique alphanumeric signature for the file, such that any modification of the content would yield a different hash value, thereby ensuring integrity verification. For long-term authenticity and proof of timing, the hash can optionally be submitted to a trusted time-stamping authority (RFC 3161) or anchored to an append-only public ledger (e.g. OpenTimestamps).
@@
  • Final disclosure. After the forecasted event has occurred, we will publish the complete set of forecast documents in an open-access repository (e.g. arXiv.org) and on the website of the Geohazards Crisis Observatory (www.geohazards-observatory.com). The corresponding hash digests will be cross-verified to demonstrate consistency with the original hashes recorded in the meta-document. A summary of all forecasts and their outcomes will also be included in the final release.
@@ 3.1 Site description
Axial Seamount is a submarine basaltic volcano located approximately 500 km offshore Oregon, USA [Chadwick2022_G3]. It has experienced three eruptive episodes over the past three decades, in 1998, 2011, and 2015. The volcano has been continuously monitored by a network of seafloor instruments [Nooner2016_Science], and since 2014 by the OOI-RCA, which includes bottom pressure recorders, tiltmeters, and seismometers [Kelley2014_MarineGeology] (see Fig. 1). This unique geophysical dataset provides an unprecedented opportunity to investigate pre-eruptive unrest processes and to test quantitative models for eruption forecasting. The 2015 eruption was previously anticipated about seven months in advance, within a one‐year window, based on pattern recognition in the geodetic time series [Nooner2016_Science]. This forecast was formulated empirically from the observed inflation pattern rather than derived from a physics‐based model. However, subsequent applications of this technique have not produced reliable forecasts due to variable inflation rates [Chadwick2024_AGU], highlighting the need for physics‐based forecasting frameworks [Acocella2024_NatRevEarthEnv]. Within the current EFE project, we aim to perform ex-ante prospective forecasting of the next eruption of this volcano using our new physics-based failure forecast model with the continuous monitoring datasets of seafloor uplift and seismicity. All datasets in this version include observations updated to 4 November 2025.
@@ Acknowledgements
This material is based upon work supported by the Ocean Observatories Initiative (OOI), a major facility fully funded by the U.S. National Science Foundation under Cooperative Agreement No. 2244833, and the Woods Hole Oceanographic Institution OOI Program Office. Q.L. is grateful for the National Academic Infrastructure for Supercomputing in Sweden (NAISS), partially funded by the Swedish Research Council through grant agreement no. 2022-06725. Q.L., P.S., and D.S. acknowledge partial support from the Norwegian Water Resources and Energy Directorate (NVE). D.S. acknowledges funding by the National Natural Science Foundation of China (Grant No. T2350710802 and U2039202), and the Center for Computational Science and Engineering at Southern University of Science and Technology. W.W.C. and S.N. acknowledge funding from US National Science Foundation awards (Grant No. OCE-2226488 and 2226445).

### v2

@@ Abstract
* Correspondence: qinghua.lei@geo.uu.seAuthor contributions. Conceptualisation: Q.L. and D.S.; Experiment design: Q.L., D.S., and W.W.C.; Code development: Q.L. and D.S.; Data acquisition: W.W.C., S.N., W.S.D.W., and M.Z.; Forecast analysis: Q.L. and D.S.; Document writing: Q.L., D.S., W.W.C., and M.Z.; Scientific discussion: all authors.
@@ 2 Experimental Design
  • Forecast release schedule. We will issue forecasts on a regular schedule announced in advance on the website of the Geohazards Crisis Observatory (www.geohazards-observatory.com). The first release was on 8 November 2025. The current document represents the second release planned on 22 December 2025. We aim to provide forecast updates once a month, or more frequently as conditions warrant.
@@
  • Digital fingerprinting. The forecast document will not be publicly released immediately. Instead, we will compute its digital fingerprint using a modern cryptographic hash function (SHA-256) under a GNU/Linux environment, in compliance with current recommendations by the National Institute of Standards and Technology. This produces a unique alphanumeric signature for the file, such that any modification of the content would yield a different hash value, thereby ensuring integrity verification (see Appendix A for a detailed description of the procedure). For long-term authenticity and proof of timing, the hash can optionally be submitted to a trusted time-stamping authority (RFC 3161) or anchored to an append-only public ledger (e.g. OpenTimestamps).
@@
  • Final disclosure. After the forecasted event has occurred, we will publish the complete set of forecast documents in an open-access repository (e.g. arXiv.org) and on the website of the Geohazards Crisis Observatory (www.geohazards-observatory.com). The corresponding hash digests will be cross-verified to demonstrate consistency with the original hashes recorded in the meta-document. A detailed description of the document authentication procedure is provided in Appendix A. A summary of all forecasts and their outcomes will be included in the final release.
@@ 3.1 Site description
Axial Seamount is a submarine basaltic volcano located approximately 500 km offshore Oregon, USA [Chadwick2022_G3]. It has experienced three eruptive episodes over the past three decades, in 1998, 2011, and 2015. The volcano has been continuously monitored by a network of seafloor instruments [Nooner2016_Science], and since 2014 by the OOI-RCA, which includes bottom pressure recorders, tiltmeters, and seismometers [Kelley2014_MarineGeology] (see Fig. 1). This unique geophysical dataset provides an unprecedented opportunity to investigate pre-eruptive unrest processes and to test quantitative models for eruption forecasting. The 2015 eruption was previously anticipated about seven months in advance, within a one‐year window, based on pattern recognition in the geodetic time series [Nooner2016_Science]. This forecast was formulated empirically from the observed inflation pattern rather than derived from a physics‐based model. However, subsequent applications of this technique have not produced reliable forecasts due to variable inflation rates [Chadwick2024_AGU], highlighting the need for physics‐based forecasting frameworks [Acocella2024_NatRevEarthEnv]. Within the current EFE project, we aim to perform ex-ante prospective forecasting of the next eruption of this volcano using our new physics-based failure forecast model with the continuous monitoring datasets of seafloor uplift and seismicity. All datasets in this version include observations updated to 19 December 2025.
@@ 4.2 Procedures
Although these three approaches differ in implementation, they are all rooted in the same LPPLS theoretical framework and provide complementary measures of forecast uncertainty. Bootstrapping captures sampling uncertainty by generating many alternative realisations of the data through residual resampling. Quantile regression probes different parts of the residual distribution to reflect asymmetric or heterogeneous fluctuations [Zhang2016_PLOSOne]. The modified profile likelihood focuses directly on how tightly the calibration can constrain the critical time $t_{c}$, propagating parametric interactions and structural uncertainty into a likelihood-based confidence measure [Filimonov2017_QuantFin]. Together, these methods illuminate distinct facets of uncertainty and yield a more complete characterisation of the predictive distribution of $t_{c}$.
@@
We stress that our forecasting methodology should be viewed as an evolving framework rather than a finalised operational tool. Significant ongoing efforts aim to learn from practice through real-time testing, refine its formulation, improve robustness, and better quantify its limitations.
@@ 5 Forecast Results
Appendix Appendix A Timestamping and Verification Procedure
To ensure full transparency and prevent any possibility of retrospective modification, each forecast in the EFE is authenticated using a simple but robust cryptographic timestamping procedure. The aim is to demonstrate unambiguously that (i) a forecast document already existed at a specific date, and (ii) its contents have not changed since that time.
For each forecast, a confidential .pdf file is created and stored privately. A SHA-256 cryptographic hash of the file is then computed using a standard GNU/Linux utility such as:
sha256sum efe_001.pdf
with the resulting 64-character hexadecimal digest serving as a unique digital fingerprint of the document. Any alteration to the file—even a single character, pixel, or whitespace—would produce a completely different hash, while reproducing the original document from the hash alone is computationally infeasible. The hash thus acts as a “digital DNA” fingerprint of the forecast document.
This hash value is recorded in the meta-document uploaded to the publicly accessible archive (e.g. arXiv.org). Because only the hash is revealed, the content of the forecast remains confidential, but its existence at the time of publication is cryptographically certified.
When all sealed forecast documents are released after the next eruption, anyone can independently verify their integrity by recomputing the SHA-256 hash of each document and comparing it with the previously published digest. Exact agreement provides definitive proof that the released file is bit-for-bit identical to the one that existed at the publicly recorded timestamp.
For convenience, we provide below the commands that readers may use to verify the authenticity of the released forecast documents. Suppose the forecast file efe_001.pdf and its previously published hash file efe_001.sha256 have both been downloaded. The integrity check is performed with:
sha256sum -c efe_001.sha256
If the file is authentic and unmodified, the terminal will report:
efe_001.pdf: OK
Alternatively, the file’s hash can be recomputed manually and compared directly with the value published in the meta-document:
sha256sum efe_001.pdf
The resulting 64-character hexadecimal digest should match the published value exactly.
If optional GPG signatures are provided (a brief cryptographic stamp created using the authors’ private key to prove authorship and authenticity), they may be verified as follows. First, import the public key included in the forecast archive:
gpg --import efe_public_key.asc
Then verify the signature associated with the hash file:
gpg --verify efe_001.sha256.asc efe_001.sha256
A message such as:
gpg: Good signature from "Qinghua Lei <qinghua.lei@geo.uu.se>"
confirms that the hash file is authentic and was signed by the forecast authors.
@@ Acknowledgements
This material is based upon work supported by the Ocean Observatories Initiative (OOI), a major facility fully funded by the U.S. National Science Foundation under Cooperative Agreement No. 2244833, and the Woods Hole Oceanographic Institution OOI Program Office. Q.L. and D.S. acknowledge partial support from the Norwegian Water Resources and Energy Directorate (NVE). D.S. acknowledges funding by the National Natural Science Foundation of China (Grant No. T2350710802 and U2039202), and the Center for Computational Science and Engineering at Southern University of Science and Technology. W.W.C. and S.N. acknowledge funding from US National Science Foundation awards (Grant No. OCE-2226488 and 2226445).